# Chiral excitation flows of multinode network based on synthetic gauge fields

### v1

@@ Appendix C The analytical solution of four-node and six-node boson-aSGF models
Hence, the dynamics can be separated into two independent subspaces, with the following Schrödinger equations and initial conditions,
@@
where the explicit form of the state ket is written as $|\psi\left(t\right)\rangle=C_{\alpha}(t)|\alpha\rangle+C_{\beta}(t)|\beta\rangle+C_{\gamma}(t)|\gamma\rangle+C_{\delta}(t)|\delta\rangle+C_{c}(t)|c\rangle$. In the decoupled $\alpha\beta$ subspace, the Hamiltonian is diagonal with energy ±$\pm 2J_{0}$, leading to the free-evolution of states $|\alpha\rangle$ and $|\beta\rangle$. However, in the $\gamma\delta c$ subspace, the zero-energy eigenstates of the four-node boson-SGF Hamiltonian $|\gamma\rangle$ and $|\delta\rangle$ interact with the auxiliary mode $|c\rangle$. After solving these time-dependent coefficients, we return to basis $\{|1\rangle,|2\rangle,|3\rangle,|4\rangle,|c\rangle\}$ and obtain the probability amplitudes for each node,
@@
Similarly, in the six-node case with $\theta_{j,j+1}=\pi/2$, we use the following unitary matrix
@@
to block-diagonalize $\mathcal{H}_{\mathrm{aSGF}}^{6}$, i.e., change the basis from $\{|1\rangle,\dots,|6\rangle,|c\rangle\}$ to $\{|\alpha\rangle,|\beta\rangle,|\gamma\rangle,|\delta\rangle,|\varepsilon\rangle,|\zeta\rangle,|c\rangle\}$, which are the eigenstates of six-node SGF Hamiltonian $\mathcal{H}_{\mathrm{SGF}}^{6}$. After the transformation, we obtain the following form,
@@
The dynamics can be likewise separated into two parts. The first part corresponds to $\alpha\beta\gamma\delta$ subspace in the upper-left $4\times 4$ block, with a diagonal matrix that indicates the simple free-evolution. The second part is restricted to the $\varepsilon\zeta c$ subspace in the lower-right $3\times 3$ block, which denotes the coupling between zero-energy states and the auxiliary mode. After solving the coefficients of the time-dependent state $|\psi\left(t\right)\rangle=C_{\alpha}(t)|\alpha\rangle+C_{\beta}(t)|\beta\rangle+C_{\gamma}(t)|\gamma\rangle+C_{\delta}(t)|\delta\rangle+C_{\varepsilon}(t)|\varepsilon\rangle+C_{\zeta}(t)|\zeta\rangle+C_{c}(t)|c\rangle$, and transforming back to the original basis $\{|1\rangle,\dots,|6\rangle,|c\rangle\}$, we obtain
@@ Appendix D Dynamics of the extended four-node network
In the extended four-node network, the energy spectrum is not equally spaced, which induces accumulated errors as the number of network nodes increases. To evaluate the imperfection of the chiral flow and investigate its origin, we here adopt a nonperturbative analysis of the system’s propagator,
@@
where $U(t,t^{\prime})$ is the time-evolution operator, and $\theta(\tau)$ denotes the Heaviside function. The propagator is related to the resolvent of the Hamiltonian $H_{\mathrm{ex}}^{N}$ by a Fourier transform
@@
$G_{+}(\omega)\equiv\lim_{\eta\to 0_{+}}G(\omega+i\eta)=\lim_{\eta\to 0_{+}}\frac{1}{E-H_{\mathrm{ex}}^{N}+i\eta}.$ (61)
Figure 11: (a) Schematics of the contour integral and the population of the 1-st node. Here, the solid line denotes analytical expression for $P_{e}^{1}$, and the dots are numerical results obtained from solving the Schrödinger equation. (b) Time evolution of the excitation on each site for $N=4$. The right region of the red dashed line represents the evolution of auxiliary emitters. (c) Normalized poles (with an $\omega_{0}$ shift) for different $N$. The unequally spaced poles result in the imperfection of fidelity.
To proceed, we divide the network hopping Hamiltonian into two parts, $H_{\mathrm{ex}}^{N}=H_{0}+V_{I}$, as depicted by the solid and dashed lines in Fig. 6(a). The unperturbed Hamiltonian $H_{0}$ contains the free Hamiltonian of all nodes as well as the hoppings between the non-corner nodes, and the interaction Hamiltonian $V_{I}$ describes the hoppings from the four corner nodes labeled as $c1$, $c2$, $c3$, and $c4$, as shown in Fig. 6(a). We constrain the dynamics within the subspace $\mathscr{E}_{0}$ formed by the corner nodes and define the projector as $P=\sum_{j=1}^{4}|c_{j}\rangle\langle c_{j}|$. Similarly, the projector of the complementary subspace is given by $Q=1-P$. Thus, the projection of $G(\omega)$ in subspace $\mathscr{E}_{0}$ is written as
@@
with $j=1,2,3,4$. Because the system preserves the chiral symmetry, the real poles of $PG(\omega)P$ are distributed symmetrically. In order to elucidate the retrieve of excitation on the 1-st node and identify the cause for the imperfection of fidelity, here we study the case of $N=4$. We find the simple poles are given by $x_{i,\pm}=\omega_{0}\pm k\sqrt{2}J_{0}$ with $i=1,2,3$ and $x_{4,\pm}=\omega_{0}\pm 2\sqrt{7}J_{0}$. By substituting the poles and calculating the sum of residues, we can then obtain the excited-state population of each node $|C_{e}^{cj}|^{2}$. The population of the 1-st node is given by
@@
As shown in Fig. 11(a), the poles of the contour integral are plotted, and the numerical result (dots) is consistent with our analytical expression (solid line). We find that each pair of poles ±$\pm x_{i}$ contributes to a cosine oscillation on the probability as in Eq. (66), and the imperfection of fidelity comes from the fact that these poles are non-uniformly distributed, i.e., the closed propagation path from $c_{1}$ to itself contains the interference with non-harmonic frequencies. Fig. 11(b) shows that the leakage mainly falls into the auxiliary emitters. We emphasize that these results are independent of the relative value of $J_{0}/\omega_{0}$ and there is no approximation made here. This implies that these hopping rates can also enter the non-perturbative regime. For larger networks, the error will accumulate as $n$ increases, along with the decreasing fidelity. This can be seen from the normalized poles $x_{i}/|x_{\mathrm{min}}|$ in Fig. 11(c), where $x_{\mathrm{min}}=\mathrm{min}\{|x_{i}|,|x_{i}|\neq 0\}$ induces the lowest oscillation frequency of $P_{e}^{j}(t)$.

### v2

@@ Appendix C Analytical solutions of four-node and six-node boson-aSGF models
Hence, the dynamics can be separated into two independent subspaces with the following Schrödinger equations and initial conditions:
@@
where the explicit form of the state ket is written as $|\psi\left(t\right)\rangle=C_{\alpha}(t)|\alpha\rangle+C_{\beta}(t)|\beta\rangle+C_{\gamma}(t)|\gamma\rangle+C_{\delta}(t)|\delta\rangle+C_{c}(t)|c\rangle$. In the decoupled $\alpha\beta$ subspace, the Hamiltonian is diagonal with energy ±$\pm 2J_{0}$, leading to the free evolution of states $|\alpha\rangle$ and $|\beta\rangle$; however, in the $\gamma\delta c$ subspace, the zero-energy eigenstates of the four-node boson-SGF Hamiltonian $|\gamma\rangle$ and $|\delta\rangle$ interact with the auxiliary mode $|c\rangle$. After solving these time-dependent coefficients, we return to basis $\{|1\rangle,|2\rangle,|3\rangle,|4\rangle,|c\rangle\}$ and obtain the probability amplitudes for each node,
@@
Similarly, in the six-node case with $\theta_{j,j+1}=\pi/2$, we use the unitary matrix
@@
to block-diagonalize $\mathcal{H}_{\mathrm{aSGF}}^{6}$, i.e., change the basis from $\{|1\rangle,\dots,|6\rangle,|c\rangle\}$ to $\{|\alpha\rangle,|\beta\rangle,|\gamma\rangle,|\delta\rangle,|\varepsilon\rangle,|\zeta\rangle,|c\rangle\}$, which are the eigenstates of six-node SGF Hamiltonian $\mathcal{H}_{\mathrm{SGF}}^{6}$. After the transformation, we obtain the following form:
@@
The dynamics can be likewise separated into two parts. The first part corresponds to the $\alpha\beta\gamma\delta$ subspace in the upper-left $4\times 4$ block, with a diagonal matrix that indicates the simple free evolution. The second part is restricted to the $\varepsilon\zeta c$ subspace in the lower-right $3\times 3$ block, which denotes the coupling between zero-energy states and the auxiliary mode. After solving the coefficients of the time-dependent state $|\psi\left(t\right)\rangle=C_{\alpha}(t)|\alpha\rangle+C_{\beta}(t)|\beta\rangle+C_{\gamma}(t)|\gamma\rangle+C_{\delta}(t)|\delta\rangle+C_{\varepsilon}(t)|\varepsilon\rangle+C_{\zeta}(t)|\zeta\rangle+C_{c}(t)|c\rangle$, and transforming back to the original basis $\{|1\rangle,\dots,|6\rangle,|c\rangle\}$, we obtain
@@ Appendix D Dynamics of the extended four-node network
In the extended four-node network, the energy spectrum is not equally spaced, which induces accumulated errors as the number of network nodes increases. To evaluate the imperfection of the chiral flow and investigate its origin, here we adopt a nonperturbative analysis of the system’s propagator,
@@
where $U(t,t^{\prime})$ is the time-evolution operator and $\theta(\tau)$ denotes the Heaviside function. The propagator is related to the resolvent of the Hamiltonian $H_{\mathrm{ex}}^{N}$ by the Fourier transform
@@
$G_{+}(\omega)\equiv\lim_{\eta\to 0_{+}}G(\omega+i\eta)=\lim_{\eta\to 0_{+}}\frac{1}{\omega-H_{\mathrm{ex}}^{N}+i\eta}.$ (58)
Figure 12: (a) Schematic of the contour integral and a plot showing the population of the first node. Here, the solid line denotes the analytical expression for $P_{e}^{1}$, and the dots are numerical results obtained by solving the Schrödinger equation. (b) Time evolution of the excitation on each site for $N=4$. The region to the right of the red dashed line represents the evolution of auxiliary nodes. (c) Normalized poles (with an $\omega_{0}$ shift) for different $N$ values. The unequally spaced poles result in the imperfection of fidelity.
To proceed, we divide the network-hopping Hamiltonian into two parts, $H_{\mathrm{ex}}^{N}=H_{0}+V_{I}$, as depicted by the solid and dashed lines in Fig. 6(a). The unperturbed Hamiltonian $H_{0}$ contains the free Hamiltonian of all nodes as well as the hoppings between the non-corner nodes, and the interaction Hamiltonian $V_{I}$ describes the hoppings from the four corner nodes labeled as $\mathscr{C}_{1}$, $\mathscr{C}_{2}$, $\mathscr{C}_{3}$, and $\mathscr{C}_{4}$, as shown in Fig. 6(a). We constrain the dynamics within the subspace $\mathscr{E}_{0}$ formed by the excited states of corner nodes and define the projector as $P=\sum_{j=1}^{4}|\mathscr{C}_{j}\rangle\langle\mathscr{C}_{j}|$. Similarly, the projector of the complementary subspace is given by $Q=1-P$. Thus, the projection of $G(\omega)$ in subspace $\mathscr{E}_{0}$ is written as
@@
Table 2: Optimal coupling parameters
number of copies coupling parameters fidelity
N=2 $2$ $\mathcal{F}=0.986$
N=3 $(2,2.26)$ $\mathcal{F}=0.960$
N=4 $(2,2.33)$ $\mathcal{F}=0.933$
N=5 $(2,2.41,2.43)$ $\mathcal{F}=0.909$
N=6 $(2,2.47,2.52)$ $\mathcal{F}=0.886$
N=7 $(2,2.51,2.59,2.67)$ $\mathcal{F}=0.865$
N=8 $(2,2.54,2.62,2.72)$ $\mathcal{F}=0.933$
N=9 $(2,2.58,2.69,2.86,2.88)$ $\mathcal{F}=0.847$
N=10 $(2,2.61,2.72,2.92,2.96)$ $\mathcal{F}=0.831$
@@
with $j=1,2,3,4$. Because the system preserves the chiral symmetry, the real poles of $PG(\omega)P$ are distributed symmetrically. To elucidate the retrieval of excitation on the first node and identify the cause of the fidelity imperfection, here we study the case $N=4$. We find that the simple poles are given by $x_{i,\pm}=\omega_{0}\pm k\sqrt{2}J_{0}$ with $i=1,2,3$ and $x_{4,\pm}=\omega_{0}\pm 2\sqrt{7}J_{0}$. By substituting the poles and calculating the sum of residues, we can then obtain the excited-state population of each node $|C_{e}^{cj}|^{2}$. The population of the first node is given by
@@
The poles of the contour integral are plotted in Fig. 12(a), and it can be seen that the numerical result (dots) is consistent with our analytical expression (solid line). We find that each pair of poles ±$\pm x_{i}$ contributes to a cosine oscillation on the probability as in Eq. (63), and the imperfection of fidelity comes from the fact that these poles are nonuniformly distributed; i.e., the closed propagation path from $\mathscr{C}_{1}$ to itself contains interference with nonharmonic frequencies. Figure 12(b) shows that the leakage mainly falls into the auxiliary nodes. We emphasize that these results are independent of the relative value of $J_{0}/\omega_{0}$ and there is no approximation made here. This implies that these hopping rates can also enter the nonperturbative regime. For larger networks, the error will accumulate as $n$ increases, along with the decreasing fidelity. This can be seen from the normalized poles $x_{i}/|x_{\mathrm{min}}|$ in Fig. 12(c), where $x_{\mathrm{min}}=\mathrm{min}\{|x_{i}|,|x_{i}|\neq 0\}$ induces the lowest oscillation frequency of $P_{e}^{j}(t)$.
Appendix E Optimized ladder network
In the unmodified ladder-network model, the fidelity of chiral flow decreases rapidly as the number of copies $N$ increases. This can be unfavorable for remote transmission between corner nodes, particularly in large networks. To improve the fidelity of chiral transmission, we transform this ladder network into a polygonal network and replace the identical coupling parameter $\beta_{c}$ between auxiliary nodes and network nodes with a set of coupling parameters $\{\beta_{j}|j=0,\cdots,\lceil N/2\rceil-1\}$. As shown in Fig. 6(d), this configuration maintains the identical treatment of corner nodes. After fixing the coupling parameters at both ends to $\beta_{0}=2$, we iterate through $\{\beta_{j}|j>0\}$ within a specific range to find the optimal parameters $\{\beta_{0}^{m}=2,\beta_{1}^{m},\cdots,\beta_{\lceil N/2\rceil-1}^{m}\}$ for each $N$, as listed in Table 2. Intuitively, the coupling strength should increase with $j$, as the auxiliary nodes that are distant from corner nodes require stronger coupling strength to carry out their role. The optimized fidelity is shown in Fig. 6(e), which demonstrates a significant improvement.
Figure 13: Scaling of the recovery time with and without the optimal parameters $\{\beta_{j}^{m}\}$. In both cases, there is a linear dependence on network size. Here, the blue solid line represents the fidelity from identical coupling, while the red dashed line denotes the results from optimal parameters.
The optimized ladder network also exhibits a shorter transmission time than the unmodified model. The circulation period of an excitation can be described by the recovery time $T_{r}$, which also characterizes the timescale for transferring an excitation between remote corner nodes. Although the recovery time shows a linear dependence on $N$, it does not increase in multiples (i.e., for a ladder network consisting of $N$ copies of the four-node aSGF model, $T_{r}<N\pi/J_{0}$). Instead, it increases much more slowly as the ladder network expands. The recovery time of the four-node aSGF model is $T_{r}=\pi/J_{0}$, while for a ladder network with $N=10$, the corresponding recovery time is less than $3\pi/J_{0}$, as illustrated in Fig. 13. When we apply the optimal parameters, there is a further reduction in $T_{r}$. As shown by the red dashed line in Fig. 13, $T_{r}$ scales sublinearly with $N$, which makes it promising for implementing remote transmission.